An electron in a computer monitor enters midway between two parallel oppositely charged plates as shown below. The initial speed of the electron before it enters the plates horizontally is 6.15 x 10 m/s. The electron is deflected a vertical distance of 4.70 mm down by the time it exits the plates. What is the magnitude and direction of the electric field between the plates? 2. electron v 6.15 x 10 m/s 4.7 mm 10 cm

if anyone can explain this in algebra physics that'd be great. can't get the right answer. thanks
Answer #1

As it can be seen, an acceleration acts in a downward direction,

Horizontal Component of motion,

V_{x}= rac{x}{t}

7 0.1 6.15 * 10

t = 1.626 * 10-9,

Consider vertical component of motion,

Using Kinematics,

s-ut 0.5at

4.7 * 10-3-0*t+0.5 * a * (1.626 * 10-9)

a = 3.555 * 1015m/s2 つつつ

The electrostatic force is responsible for this acceleration

F_{e}= ma

q*E= ma

ma

(9.1 * 10-31) * (3.555 1015) 1.6 10-19

E-2.02 19 * 104 Лус (downwards)